OpenStudy (anonymous):

can you solve the equation 3/4 (6x + 9 ) = 14 by using the division property of equality? explain

OpenStudy (anonymous):

sure can but can you get started?

OpenStudy (anonymous):

im not sure, first time seeing a question like this.

OpenStudy (anonymous):

ok can you distribute ?

OpenStudy (anonymous):

just have to explain how it can be solved. can you distribute with a fraction?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

go ahead and do that

OpenStudy (anonymous):

.75(6x) + .75(9) ?

OpenStudy (anonymous):

yes or you can write it this way \[\frac{3}{4}*6x+\frac{3}{4}*9\] now can you simplify this \[\frac{3}{4}*6x+\frac{3}{4}*9=14\]

OpenStudy (anonymous):

4.5x + 6.75 = 14 4.5x + 6.75 - 6.75 = 14 - 6.75 4.5x = 7.25 ?

OpenStudy (aripotta):

yea, then divide both sides by 4.5

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

@AriPotta give him a chance

OpenStudy (aripotta):

what do you mean? :l i'm just telling him the next step

OpenStudy (anonymous):

well, the lasts tep is always simplification, so :P

OpenStudy (anonymous):

4.5x = 7.25 now divide by 4.5 what do you get?

OpenStudy (anonymous):

1.6

OpenStudy (aripotta):

personally, i'd change everything back into a fraction. because your answer as a decimal would be repeating

OpenStudy (anonymous):

we have to show it like that, its very annoying.

OpenStudy (aripotta):

oh, it has to be a decimal?

OpenStudy (anonymous):

if its done any other way, it is counted wrong, even if the answer is correct

OpenStudy (aripotta):

wow, that's stupid.

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

(6x + 9 ) = 14 \[\frac{3}{4}6x + \frac{3}{4}9=14\] \[\frac{18}{4}x + \frac{27}{4}=14\] \[\frac{18}{4}x =14- \frac{27}{4}\] \[\frac{18}{4}x =\frac{29}{4}\] \[x =\frac{29}{4}*\frac{4}{18}\] \[x =\frac{29}{18}\]